OpenStudy (anonymous):

How do you rationalize the denominator in 6 over the square root of 2? Please show step by step and express answer in lowest radical form with no decimal.

OpenStudy (lalaly):

\[\frac{6}{\sqrt2} \times \frac{\sqrt2}{\sqrt2}\]\[=\frac{6\sqrt2}{2} = 3\sqrt2\]

OpenStudy (anonymous):